Western Airlines wants to design a hub system in the United States. Each hub is used for connecting flights to and from cities within 1500 miles of the hub. Western runs flights among the following cities: Atlanta, Boston, Chicago, Denver, Houston, Los Angeles, New Orleans, New York, Pittsburgh, Salt Lake City, San Francisco, and Seattle. The company wants to determine the smallest number of hubs it needs to cover all these cities. A city is covered if it is within 1500 miles of at least one hub. The table on the left lists actual distances between cities.

Find the location of the hubs.Input data Mile limit Background: Western Airlines wants to design a hub system in the United States. Each hub is used for coWhere are the hub(s) located?
Answer #1

Solution: An integer programming problem has been formulated and solved with Excel solver. Following are the snapshots of Excel model and solver window:

Total_Hubs - X fx Western Airlines Hubs Planning Solver Parameters A Western Airlines Hubs Planning Set Objective: Total_Hubs

X f r Total_Hubs - A Solver Parameters B 0 Western Airlines Hubs Planning C D E F G H 0 0 0 1 0 I 0 J 0 K 1 L 1 M 1 SE Set Ob

Following snapshots show the formulas used in various cells:

F G H I J K L M E Western Airlines Hubs Planning wN AT сн O 3 Input Data Distance between cities AT BO CH DE 0 HO LA NO 1037

E F G H I T K L M - IF(E16<=1 IF(F16<=1 -IF(G16<1 -IF(H16<1 -IF(116<=1 =1F(J16<=1 IF(K16<=1 -IF(L16<=1 IF(M16 SE -IF(B16<=150

Answer: As we see in the optimized state of the model (in the snapshots), the recommended hub locations are Atlanta and Denver.